OpenStudy (anonymous):

Find the vertex of the parabola whose equation is y = -2x^2 + 8x - 5. (2, 3) (2, 19) (2, 27)

OpenStudy (campbell_st):

an easy method for a quadratic \[ax^2 + bx + c \] is to find the line of symmetry... \[x = \frac{-b}{2a}\] and the vertex is on the line... then substitute this value into the equation to find the corresponding y value... and then you have the vertex in your question a = -2 and b = 8 hope it helps

OpenStudy (anonymous):

I know that, and I did it, but the answer I got (even after repeating everything numerous times) was the same and apparently it's not right.

OpenStudy (campbell_st):

well what did you get for the line of symmetry x = ?

OpenStudy (anonymous):

The answer I got is (2, 15) but obviously that isn't correct lol

OpenStudy (campbell_st):

well the x value is correct... substituting y = -2(2)^2 + 8(2) - 5 y = -8 + 16 - 5 so y = 3 thats my best guess

OpenStudy (anonymous):

Thank you.. I'll just go with that then lol

OpenStudy (campbell_st):

the x value is correct... the error was just in the substitution and evaluation...

OpenStudy (whpalmer4):

You should figure out why you're making a mistake evaluating \[y = -2(2)^2+8(2)-5\]so that you can avoid that mistake in the future...

OpenStudy (anonymous):

I'll deal with the future when it gets here! This question is closed, thanks for the help.